a.      Let A and B and C be sets, prove that A∩(BC) = (A∩B) ( A∩C). 

b. a.      The number x is multiplied by itself and the result y is less than the number x.

For example, x*x = x2 = y and y < x. 

i.   State a set name in which numbers behave in such an assertion above.

ii.     State any number in the stated set in (a) above and verify the assertion. 


1
Expert's answer
2021-03-09T07:51:15-0500

a. Distributive law "A\cap(B\cup C)=(A\cap B)\cup (A\cap C)"

Suppose that "x\in A\cap (B\cup C)." Then "x\in A" and "x\in B\cup C." By the definition of union, it follows that "x\in A" and "x\in B" or "x\in C" (or both).

In other words, we know that the compound proposition "(x\in A) \land((x\in B)\lor(x\in C))" is true.

By the distributive law for conjunction over disjunction, it follows that "((x\in A) \land(x\in B))\lor((x\in A) \land(x\in C))."

We conclude that either "x\in A" and "x\in B" or "x\in A" and "x\in C" .

By the definition of intersection, it follows that "x\in A\cap B" or "x\in A\cap C."

Using the definition of union, we conclude that "x\in (A\cap B)\cup (A\cap C)." We conclude that "A\cap(B\cup C)\subseteq(A\cap B)\cup (A\cap C)."


Now suppose that "x\in A\cap(B\cup C)." Then, by the definition of union, "x\in A\cap B" or "x\in A\cap C."

By the definition of intersection, it follows that "x\in A" and "x\in B" or that "x\in A" and "x\in C."

From this we see that "x\in A," and "x\in B" or "x\in C." Consequently, by the definition of union we see that "x\in A" and "x\in B\cup C." Furthermore, by the definition of intersection, it follows that "x\in A\cap (B\cup C)."

We conclude that "(A\cap B)\cup (A\cap C)\subseteq A\cap(B\cup C)."

This completes the proof of the identity.


The membership table for these combinations of sets is shown in Table.


"\begin{matrix}
 A & B & C & B\cup C & A\cap (B\cup C) \\
 1 & 1 & 1 & 1 & 1 \\
1 & 1 & 0 & 1 & 1 \\
1 & 0 & 1 & 1 & 1 \\
1 & 0 & 0 & 0 & 0 \\
0 & 1 & 1 & 1 & 0 \\
0 & 1 & 0 & 1 & 0 \\
0 & 0 & 1 & 1 & 0 \\
0 & 0 & 0 & 0 & 0 \\
\end{matrix}"


"\begin{matrix}
 A & B & C & A\cap B & A\cap C & (A\cap B)\cup (A\cap C) \\
 1 & 1 & 1 & 1 & 1 & 1 \\
1 & 1 & 0 & 1 & 0 & 1 \\
1 & 0 & 1 & 0 & 1 & 1 \\
1 & 0 & 0 & 0 & 0 & 0 \\
0 & 1 & 1 & 0 & 0 & 0 \\
0 & 1 & 0 & 0 & 0 & 0 \\
0 & 0 & 1 & 0 & 0 & 0 \\
0 & 0 & 0 & 0 & 0 & 0 \\
\end{matrix}"

Because the columns for "A\cap(B\cup C)" and "(A\cap B)\cup (A\cap C)" are the same, the identity is valid.


b. "y=x^2<x"

i.


"x^2<x=>x^2-x<0=>x(x-1)<0"

"=>x\in(0, 1)"

The opened interval "(0, 1)."


ii. Let "x=\dfrac{1}{2}." Then


"y(\dfrac{1}{2})=(\dfrac{1}{2})^2=\dfrac{1}{4}<\dfrac{1}{2}=x"